Two balls are drawn in succession without replacement from the urn containing 5 yellow balls and 6 green balls. Let Z be the random variable representing the number of green balls. Construct a probability distribution of Z in tabular, graphical, and equation form.


1
Expert's answer
2021-06-21T11:21:54-0400

Given,

5 yellow balls and 6 green balls.

Total=11


Z is a random variable So value of Z can be 0,1 or 2 as two balls are drawn in succession.


When there is 0 green balls SO both balls are yellow.


"P(Z=0)=\dfrac{^5C_2}{^{11}C_2}=0.18"



Now If there is one green ball ,This can be done in two ways either green first then yellow or vice versa.


"P(Z=1)=\dfrac{5}{11}\times \dfrac{6}{10}+\dfrac{6}{11}\times \dfrac{5}{10}=0.545"


When both balls are green.


"P(Z=2)=\dfrac{^6C_2}{^{11}C_2}=0.272"


Probability Distribution table is-





Graphically PD is-



Equation form of PD is-


"P(x)=0.09x^2+0.45x+0.18"